OpenStudy (anonymous):

The amount of revenue a company makes per day by selling x items is given by the function f(x) = 15x – 0.1x2. How many items should be sold if the company wants to maximize their profit?

OpenStudy (saifoo.khan):

\[-\frac{b}{2a} \]\[-\frac{15}{-0.1}\]\[150\]

OpenStudy (anonymous):

\[-\frac{15}{-0.2}\] 75

OpenStudy (anonymous):

Thanks guys ;)

OpenStudy (saifoo.khan):

Ops. My bad. lol